OpenStudy (anonymous):

Indices, Surd and Numeral Systems #1 Question 40bii (modified) How to convert a binary number into hexadecimal number without converting it into decimal number? Eg: \(1111001110_2\)

OpenStudy (amistre64):

1 hex digit takes 4 binary digits

OpenStudy (anonymous):

Yes. I understand that. But what should I do next?

OpenStudy (amistre64):

0011 1100 1110 3 C E

OpenStudy (phi):

group into sets of 4 change each set into a hex number

OpenStudy (amistre64):

im sure theres a mathical thing that can happen; but they only really differ from 10 to 15

OpenStudy (anonymous):

How do you find : 0011 ?

OpenStudy (amistre64):

0011 = 3 in binary; so it equals 3 in hex

OpenStudy (amistre64):

from 0 to 9 hex and decimals are the same; they differ at 10 - 15 (A - F)

OpenStudy (amistre64):

splitting the binary into groups of 4, starting from the right side, is helpful

OpenStudy (anonymous):

So, in my example 1111001110 First, group them with 4 digits, counted from the right to the left 11 1100 1110 Since there are insufficient number for the first group, add enough zeros to make it a group of 4. Then it becomes 0011 1100 1110 For the first group, 0011 = 2^2 + 1 = 3 <- first digit of the hex. number For the second group, 1100 = 2^3 + 2^2 = 12 = C <- second digit in the hex. number For the third group, 1110 = 2^3 + 2^2 +2 = 14 = E <- last digit in the hex. number So, the final answer is \(1111001110_2 = 3CE_{16}\) Is this the case?

OpenStudy (amistre64):

thats the case

OpenStudy (anonymous):

Thanks!